### v2

# AfriMed-QA: A Pan-African, Multi-Specialty, Medical Question-Answering Benchmark Dataset

Abstract
Recent advancements in large language model (LLM) performance on medical multiple-choice question (MCQ) benchmarks have stimulated interest from healthcare providers and patients globally. Particularly in low-and-middle-income countries (LMICs) facing acute physician shortages and lack of specialists, LLMs offer a potentially scalable pathway to enhance healthcare access and reduce costs. However, their effectiveness in the Global South, especially across the African continent, remains to be established. In this work, we introduce AfriMed-QA 22footnotetext: Our dataset is available at:
https://huggingface.co/datasets/intronhealth/afrimedqa_v2
, the first large-scale Pan-African English multi-specialty medical Question-Answering (QA) dataset, 15,000 questions (open and closed-ended) sourced from over 60 medical schools across 16 countries, covering 32 medical specialties. We further evaluate 30 LLMs across multiple axes including correctness and demographic bias. Our findings show significant performance variation across specialties and geographies, MCQ performance clearly lags USMLE (MedQA). We find that biomedical LLMs underperform general models and smaller edge-friendly LLMs struggle to achieve a passing score. Interestingly, human evaluations show a consistent consumer preference for LLM answers and explanations when compared with clinician answers.
AfriMed-QA: A Pan-African, Multi-Specialty, Medical Question-Answering Benchmark Dataset
Tobi Olatunji1,2,4, Charles Nimo2, Abraham Owodunni1,3, Tassallah Abdullahi4,12, Emmanuel Ayodele1, Mardhiyah Sanni1,4, Chinemelu Aka1, Folafunmi Omofoye4, Foutse Yuehgoh4, Timothy Faniran4, Bonaventure F. P. Dossou4,10, Moshood Yekini4, Jonas Kemp6, Katherine Heller6, Jude Chidubem Omeke4, Chidi Asuzu MD4, Naome A. Etori4,11, Aimérou Ndiaye5, Ifeoma Okoh5, Evans Doe Ocansey5, Wendy Kinara7, Michael Best2, Irfan Essa2,6, Stephen Edward Moore8, Chris Fourie9, Mercy Nyamewaa Asiedu6♠ 1Intron, 2Georgia Institute of Technology, 3The Ohio State University, 4BioRAMP, 5Masakhane, 6Google Research, 7Kenyatta University, 8University of Cape Coast, 9SisonkeBiotik, 10MILA Quebec, 11University of Minnesota, 12Brown University tobi@intron.io
11footnotetext: Authors marked with $\spadesuit$ are senior authors.
1 Introduction
Figure 1: Methodology Overview
Table 1: Comparative analysis of the AfriMed-QA dataset with health datasets.
Feature AfriMed-QA BioASQ Krithara et al. (2023) MedQA-SWE Hertzberg and Lokrantz (2024) MedQA Jin et al. (2020) MedMCQA Pal et al. (2022b) PubMedQA Jin et al. (2019) MMLU Hendrycks et al. (2020) HealthSearchQA Singhal et al. (2022a)
Dataset Size 15,275 4,721 3,180 12,723 193,155 1,000 15,908 3,173
Question Types MCQ, SAQ, yes/no, Factoid, MCQ MCQ MCQ Yes/No/maybe, MCQ Consumer Queries
Consumer Queries List, Summary Factoid, List
Clinical Scenarios $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
Answer Options 2-5 (MCQs) $\times$ 5 (MCQs) 4 (MCQs) 4 (MCQs) $\times$ 4 (MCQs) $\times$
Correct Answers $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Answer Rationale $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
Question Source $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
Large language models (LLMs) have gained popularity in specialized domains such as finance Wu et al. (2023b), medicine Singhal et al. (2022b), and climate Thulke et al. (2024). In the medical field, LLMs like Med-PaLM Singhal et al. (2022b), PMC-LLaMA Wu et al. (2023a), and GPT-4 Achiam et al. (2023) have shown impressive performance in tasks such as summarizing clinical notes and answering medical questions with high accuracy Liu et al. (2024c); Eriksen et al. (2023); Abdullahi et al. (2024); Liu et al. (2024a). Especially in low-resource settings, these models have the potential to improve clinician productivity, accessibility, operational efficiency, and enable multilingual clinical decision support Yang et al. (2023); Gangavarapu (2023).
Despite their success on existing medical benchmarks, it is uncertain whether these models generalize to tasks involving linguistic variations, even within English, localized cultural contexts, and region-specific medical knowledge, highlighting the need for more diverse benchmark datasets. Current evaluations rely on publicly available digital resources Kung et al. (2023); Jin et al. (2019, 2021), but these may not translate to out-of-distribution datasets, such as those from African countries.
To address this gap, we introduce AfriMed-QA, a dataset of 15,275 English, clinically diverse questions and answers, 4,000+ expert multiple-choice questions (MCQs) with answers, over 1,200 open-ended short answer (SAQs) with long-form answers, and 10,000 consumer queries (CQ), to rigorously assess LLM performance for correctness and demographic bias. We evaluate 30 large, small, open, closed, biomedical, and general LLMs using quantitative and qualitative approaches. While the development of the dataset is still in progress, this work establishes a foundation for acquiring diverse and representative health benchmark datasets across LMICs.
2 Related Work
2.1 Medical Domain Benchmark Datasets for LLMs
The Open Medical LLM Leaderboard Pal et al. (2024) tracks, ranks, and evaluates LLM performance on medical question-answering tasks across diverse datasets, including MedQA (USMLE) Jin et al. (2021), PubMedQA Jin et al. (2019), MedMCQA Pal et al. (2022a), and subsets of MMLU Hendrycks et al. (2020). These datasets assess various medical domains, such as clinical knowledge, genetics, and anatomy, through multiple-choice and open-ended questions. CMExam Liu et al. (2023) from the Chinese National Medical Licensing Examination offers MCQs with detailed annotations, while Q-Pain Log’e et al. (2021), Medication QA Abacha et al. (2019), LiveQA Liu et al. (2020), MultiMedQA Singhal et al. (2022a), and EquityMedQA Pfohl et al. (2024) cover various medical QA challenges. Table 1 compares AfriMed-QA to other medical QA benchmarks.
2.2 Evaluating LLMs for Health-Specific Tasks
There have been several studies evaluating LLMs for medical standardized exams and for various clinical tasks using approaches like zero-shot, finetuning, developing benchmarking metrics and running human evalutations Jahan et al. (2024); Singhal et al. (2022b); Fleming et al. (2024); Umapathi et al. (2023); Chen et al. (2024). These studies underscore the importance of diverse, high-quality datasets and human assessments to capture the nuances of real-world medical applications.
Reddy et al. Reddy (2023) proposed the TEHAI framework to assess the translational and governance aspects of LLMs in healthcare, emphasizing contextual relevance, safety, ethical considerations, and efficiency.
Our evaluation extends TEHAI and work from Singhal et al. Singhal et al. (2022b), and Pfohl et al. Pfohl et al. (2024) by incorporating specialty and region-specific dimensions, ensuring a comprehensive assessment of LLMs. Metrics such as local expertise, harmfulness, and bias are included to cover ethical dimensions and governance. By addressing correctness, omission, hallucination, and reasonableness, we thoroughly evaluate AI systems in healthcare.
The AfriMed-QA dataset aims to (1) integrate geo-culturally diverse datasets, specifically those from African LMICs that have historically relied on paper-based records, and local health data and are underrepresented in LLM training and evaluation; and (2) expand healthcare LLM benchmark datasets to include African consumer/patient-based queries. This enables LLM training and evaluation on a broad spectrum of medical data, creating more robust, inclusive, and practical AI solutions for Africa-centric applications.
3 AfriMed-QA Dataset
We introduce AfriMed-QA, the first large-scale Pan-African multispecialty medical Question-Answer dataset designed to evaluate and develop equitable and effective LLMs for African healthcare. Figure 1 details our methodology and data collection process.
Question Tier Type Count Total
Expert MCQ 3,910 4,269
SAQ 359
Crowdsourced MCQ 129 11,006
SAQ 877
CQ 10,000
Total 15,275
Table 2: Dataset statistics
Table 2 and 9 shows dataset statistics. Multiple-choice (MCQ) professional medical exam questions include a question, 2-5 alternative answer options, the correct answer(s), and the rationale for the correct answer (answer option count distribution is shown in Appendix Tab 8). Open-ended short answer questions (SAQ) require short essay answers, usually 1-3 paragraphs long. Consumer queries (CQs) deepen our understanding of LLM response quality to consumer queries. To maximize the diversity of consumer questions, we leveraged a curated list of 472 medical conditions, symptoms, and patient complaints common in African communities across 32 specialties, to create culturally appropriate prompts that help elicit diverse questions from clinical and non-clinical crowd workers. Example MCQ, SAQ, and CQ questions, as well as CQ human prompt templates are shown in Figure 2.
This dataset was crowd-sourced from 621 contributors (Female 55.56%, Male 44.44%) from over 60 medical schools across 16 countries (Table 9), covering 32 medical specialties including Obstetrics & Gynecology, Neurosurgery, Internal Medicine, Emergency Medicine, Medical genetics, Infectious Disease, and others. Appendix Table 11 shows the Specialty distribution. Human answers and explanations are provided for 5,444 questions. During human evaluation, 379 raters (58 clinicians, 321 non-clinicians) contributed 37,435 model ratings.
Figure 2: Question Samples for multiple choice questions (MCQs), Short Answer Questions (SAQs), and Consumer Queries
3.1 Data Collection
We adapted a web-based platform previously developed by Intron Health  [^1] Intron Health’s biomedical crowd-sourcing platform https://speech.intron.health to crowd-source accented and multilingual clinical speech data at scale across Africa Olatunji et al. (2023). Custom user interfaces were developed to collect each question type, for quality reviews, and for blind human evaluation of LLM responses. The MCQ User Interface and other details about the data collection tool can be found in Appendix Figure 13.
3.2 Contributor Recruitment and Instructions for Data Collection and Evaluation
Contributors: Medical trainees and clinicians were recruited to contribute questions through referrals from existing Intron Health’s web-based platform contributors, medical associations, and via social media. Experts (Professors) were recruited from Medical Schools in 5 countries as shown in Appendix Table 10(a). Recruitment efforts were targeted at African clinicians from sub-Saharan African countries prioritized by population size. To maximize geographic representation, each contributor was limited to a maximum of 300 questions and answers. Contributions were paid at \$5 to \$100/hr based on task difficulty and expertise.
Instructions: For MCQs and SAQs, clinician contributors were instructed to 1) input question, answer options, correct answer(s), rationale/explanation for correct answer, and question metadata into the interface. Experts provided MCQ answers with no explanations. 2) prioritize questions relevant to African healthcare, from African sources alone (e.g. no USMLE prep questions). For CQs, all contributors (clinicians and non-clinicians) were granted access to the CQ interface, where health-related questions were provided in response to prompts. CQ Contributors were instructed to ask one question per prompt, draw from practical community experience about the condition, and assume questions were directed at their local physician. Clinician contributors were then granted privileged access to a dedicated interface to provide human answers to consumer queries.
Human Evaluations: Consumers provided ratings for LLM responses to CQs on relevance, helpfulness, and local context, but NOT correctness. Due of the higher expertise required, only confirmed clinicians were granted access to projects rating MCQ, SAQ, and CQ answers. Clinician status was confirmed through submitted credentials and background checks reviewing publicly available information about them. Raters were randomly assigned (double-blind) to the answer source (human or LLM). More details on human evaluations are provided in Section 4.3
Quality Review: We utilized the contributor quality review process described in Olatunji et al. (2023). Contributors were rigorously evaluated by a team of clinicians. Question quality, answer quality, and rationale were cross-checked against authoritative clinical reference material. Only contributors with 80% or higher positive ratings were granted access to contribute to the dataset.
4 Approach
4.1 LLM Selection
We evaluate 30 LLMs (Table 3) including open-source and proprietary LLMs, general-purpose and biomedical LLMs, mixture-of-experts, and models of varying sizes from 3B to over 540B parameters Chowdhery et al. (2023)). Of these, 13 are proprietary while 17 are open-source. We evaluate LLMs like Phi-3 Abdin et al. (2024), GPT-4 Achiam et al. (2023), MedLM, Claude 3 Anthropic (2023), OpenBioLLM Pal and Sankarasubbu (2024), Gemini Team et al. (2023), Meditron Chen et al. (2023), and MedLlama Wu et al. (2023a).
4.2 Quantitative Evaluation
For multiple-choice tasks, accuracy is measured by comparing LLM’s single-letter answer choice [A,B,C,D,E] with the reference. For open-ended questions, semantic similarity is measured using BertScore, comparing the generated response from the language model against a reference answer Zhang* et al. (2020).
Figure 3: AfriMedQA: Expert-MCQ accuracy, SAQ, and CQ Bertscore
(a) Top: MedQA vs AfriMedQA MCQ Accuracy. Bottom: Effect of Explanations
(b) MCQ accuracy by specialty
Figure 4: Breakdown of MCQ accuracy by specialty and dataset
Figure 5: MCQ accuracy by country
4.3 Qualitative: Human Evaluations and Evaluation Axes
LLM responses to a fixed subset of questions (n=3000, randomly sampled) were sent out for human evaluation on the Intron Health crowd-sourcing platform. Adapting the evaluation axes in Singhal et al. (2023), we collected human evaluations in two categories: (1) Non-clinicians were instructed to provide ratings to CQ LLM responses to determine if answers were relevant, helpful, and localized; (2) Clinicians were instructed to provide ratings to the LLM’s MCQ, SAQ, and CQ responses to determine if answers were correct and localized, if omissions or hallucinations were present, and if potential for harm exists. Evaluation axes and exact instructions are detailed in Appendix section A.1.
Ratings were on a 5-point scale representing the extent to which the criteria were met. 1 represents “No" or “completely absent", and 5 represents “Yes" or “Absolutely present". Raters were blinded to the answer source (model name or human). Each rater evaluated multiple LLMs or human answers in random blind sequence.
4.4 Experiment Setup
Checkpoints for open-source models were sourced from HuggingFace and Google’s Vertex AI Studio, while proprietary models were accessed through developers’ API using default hyperparameters. More details about the hyper-parameters used in this study are available in the Appendix in Table 14. Aligning with our community participatory design approach, experiments by multiple collaborators ran on various hardware types including NVIDIA L4, NVIDIA T4, and A100.
5 Results
5.1 Benchmark evaluations
AfriMed-QA MCQ, SAQ, and CQ evaluation results are shown in Figure 3. Accuracy ranges from 0.17 (Gemma-2B, the smallest LLM in this study) to 0.79 (GPT-4o). GPT-4o, Claude-3.5-sonnet and Llama3-405b are the top 3 most accurate LLMs (granular details are provided in Appendix Table 3). Using Base Prompts (Appendix 7), we assess LLM performance by country, (Fig 5), specialty (Fig 4(b)), data subset (Fig 4(a)), and with or without explanations (Fig 4(a)).
5.2 Human evaluations
Figure 6 shows results from clinician and non-clinician human evaluators on the dataset across various axes for LLM and Human responses. Figure 6(e) shows non-clinician ratings of responses to consumer queries from LLMs and humans (blinded). Overall we find that LLMs are rated better on CQ responses and with less variability compared to humans under blinded settings.
(a) Correctness
(b) Irrelevant information
(c) Omission of relevant information
(d) Possibility of harm
(e) Consumer ratings of CQs
Figure 6: Clinician and consumer blind evaluations of human and LLM answers showing mean ratings and confidence intervals across various axes.
6 Discussion
The experimental results of our study on AfriMed-QA reveal several key insights and trends.
6.1 The dominance of large models may be unfavorable for low-resource settings
Figure 3 shows a wide range of LLM accuracies on Expert MCQs with the best models (large, 100B-2T, proprietary) scoring over 75%, smaller models (under 10B) clustering in the 40-60% range, and medium models (11B-70B) scoring between 60 and 75% (details in Appendix Table 3). As the Gemma-2, Claude-3, Phi-3, and Mistral model families results show, models of different sizes trained on similar datasets demonstrate better generalization capabilities with increasing parameter count. Mixture-of-Expert (MoE) models Mistral-8x7B outperforms its biomedical and general 7B variants by a wide-margin further confirming the correlation between model capacity and performance. Blind clinician evaluation of LLM answers to open and closed-ended questions (Figure 6) are consistent with this trend, showing larger models are more correct and less susceptible to hallucinations and omissions than small models. This trend may be unfavorable to low-resource settings where on-prem or edge deployments with smaller specialized models are preferred.
6.2 Localization is still a challenge
Figure 4(a) (top left) reveals an unmistakable performance gap between USMLE (MedQA) (orange line) and AfriMed-QA Expert MCQs (blue line), with proprietary GPT-4o, Claude-3.5-sonnet and Gemma-2B showing an 8.86, 5.57, and 15.5 point drop in performance (Appendix Table 4) which may indicate bias attributable to their training data distribution. Figure 4(a) further shows differences in question difficulty by source. Since MCQs that mention African cities or locations (e.g. A 50-year old patient with recent travel to Uganda…) were mostly contributed by Trainees (green line), relatively higher LLM accuracy on this subset (for small and large LLMs) suggests they were relatively easier for LLMs when compared with Expert MCQs (blue line) and MCQs that did not mention African geographies (red line). This finding is counter-intuitive and requires further investigation.
6.3 Evidence of progress in LLM reasoning abilities
Performance of the GPT model series (3.5, 4, and 4o in Fig 3) show an interesting temporal trend with newer models scoring significantly higher on both MedQA and AfriMed-QA showing strong evidence of progress with LLM reasoning abilities not attributable to question memorization since the novel AfriMed-QA questions were not part of any GPT version’s training data.
6.4 Domain-specific biomedical LLMs still struggle
Figure 3 shows general models outperform and generalize better than biomedical models of similar size (8B and 70B). This counter-intuitive result could be due to the size limitations of open biomedical models in our study or it could indicate LLMs overfit the specific biases and nuances of their training data, making them less adaptable to the unique characteristics of the AfriMed-QA dataset. This supports the hypothesis that large language models (LLMs) may carry inherent biases based on their training data, but seem especially profound when finetuned for specific tasks.
6.5 Specialties must select LLMs with caution
Figure 4(b) shows a clear top-down, left-right trend indicating which LLMs are more reliable in certain specialties. Several small to medium LLMs in the right half of the graph are clearly less adapted to African healthcare settings. While larger closed LLMs seem to generalize across specialties, our results reveal that LLMs perform better on Medical Specialties (Rheumatology, Nephrology, Gastroenterology, Endocrinology, Pulmonary, etc) when compared with other specialties like Surgery, Pathology, Pediatrics, Infectious Diseases, and Obstetrics & Gynecology that are very important in LMICs, a striking trend that requires further investigation.
6.6 Performance variation by country
Figure 5 shows a clear difference in difficulty of expert question from South Africa and Nigeria for LLMs. This requires further investigation but could be a result of differences in specialty distributions of expert MCQs per country. For example, expert questions from South Africa are dominated by Pediatrics, a difficult specialty for LLMs as shown in Figure 4(b) while a significant number of Pathology and Obstetrics and Gynecology questions come from Nigeria. Country-Specialty counts are detailed in Appendix Tables 7, 6, and 10.
6.7 Positive and negative insights from LLM explanations
We investigated the effect of generating explanations on LLM accuracy (Fig 4(a) bottom left) and found that, contrary to the general notion that model explanations improve LLM accuracy Wei et al. (2022), in the context of MCQs, post-processing (regex or pattern matching) challenges with automatically extracting the answer option selected from the explanation led to subpar results and LLMs scored higher without explanations (Appendix 12). For example, Claude Opus generally struggled with pattern consistency, producing variations like “The most appropriate … <option>" or “The doctor should respond … <option>" instead of simply providing its answer– “Option B" followed by its rationale. This highlights challenges with the ability of LLMs to produce consistent or structured outputs in response to instructions Liu et al. (2024b).
6.8 Consumers prefer LLM answers
Consumer and clinician human evaluation of LLM answers to CQs (Fig 6(e)) revealed an overwhelming preference for LLM responses as they were consistently rated to be more complete, informative, and relevant when compared with clinican answer brevity (Fig 2). Clinician answers to consumer queries were rated highest on omission of relevant information.
6.9 Potential for harm, omissions, and hallucinations still persist
Figure 6(e) showed that smaller open general and biomedical LLMs like Llama-3-8b and JSL-Med-llama-8b had the highest count of answers with hallucinations, omissions, and the potential for harm in MCQ, open-ended SAQ, and CQ answers. Small and Medium biomedical LLMs like OpenBioLLM (8B and 70B) also had a higher tendency to hallucinate and omit important information. Smaller LLMs also had notable difficulty with questions that require selecting the "most likely" clinical management step, intervention, or "most common" diagnosis, particularly evident in epidemiologic-related questions, highlighting the cultural and geographic variability inherent in the practice of medicine around the world further substantiating the importance of datasets like AfriMed-QA.
7 Conclusion
In this work, we introduce AfriMed-QA, the first large-scale multi-specialty Pan-African medical Question-Answer (QA) dataset comprised of 15k MCQs, SAQs and CQs, designed to evaluate and develop equitable and effective LLMs for African healthcare. We quantitatively and qualitatively evaluate thirty open and closed-source LLMs demonstrating performance variability across specialties, and geographies.
8 Limitations and Future Work
Although this is the first large-scale, multi-specialty, indigenously sourced Pan-African dataset of its kind, it is by no means complete. Over 60% of the expert MCQ questions came from West Africa. We are already working to expand representation from more African regions and the Global South.
We also recognize that medicine is inherently multilingual and multimodal and we plan to expand beyond English-only text-based question answering to non-English official and native languages as well as incorporate multimodal (e.g. visual and audio) question answering. Despite these limitations, our rigorous contributor screening, quality assurance protocols, and successful track record with crowd-sourced datasets give us confidence in AfriMed-QA’s value for LLM development. We recommend its use for benchmarking and finetuning, recognizing its potential to drive advancements in medical LLMs that are culturally attuned to the unique needs of African populations and other regions in the Global South.
Acknowledgements
Support for this project was provided by Google Research, PATH, and the Bill & Melinda Gates Foundation [Gates INV-068056]. We appreciate the excellent execution by the Intron Health Data Team who managed this pan-African data collection effort including recruiting contributors, developing custom UIs, quality reviews, and multi-currency contributor payments. We appreciate the invaluable support provided by the BioRAMP researchers, whose collaboration and insights have been fundamental to our research. We also extend our thanks to the Masakhane Community for their expert contributions, the Sisonke Biotik community for their enthusiastic participation and commitment, and the Federation of African Medical Students’ Associations [FAMSA] for their relentless recruitment and community engagement. These partnerships have been instrumental in enhancing the scope and impact of our work, and we appreciate their dedication to advancing knowledge in our field.
Appendix A Appendix
Table 3: Zero-Shot Expert Evaluation of Large Language Models: Accuracy and BERT Scores on Multiple Choice Questions, Short Answer Questions, and Consumer Queries. Some model names are shortened for brevity.
Model MCQ-Expert SAQ Bertscore CQ Bertscore Domain Access Size Training
Gpt-4 0.7568 0.8727 0.8385 General Closed 1.8T Instruct
Gpt-4o 0.7928 0.8825 0.8254 General Closed 12B Instruct
Gpt-3.5 Turbo 0.5629 0.8813 0.0963 General Closed 175B Instruct
Gpt-4o mini 0.7176 0.8808 - General Closed $\sim$8B Instruct
Claude-3.5 Sonnet 0.777 0.8574 - General Closed - Instruct
Claude-3 Sonnet 0.6504 0.8719 0.8141 General Closed 70B Instruct
Claude-3 Opus 0.7455 0.8696 0.8172 General Closed 2T Instruct
Claude 3 Haiku 0.6639 0.8656 0.8163 General Closed 2T Instruct
Gemini Pro 0.6036 0.8601 0.8213 General Closed 540B Instruct
MedLM 0.6036 0.8633 0.8303 Biomedical Closed - Instruct
Gemini Ultra 0.739 0.8716 0.8362 General Closed 1.56T Instruct
MedPalm2 - 0.8716 0.8379 Biomedical Closed 540B Instruct
llama3-405b-instruct-maas 0.7627 - General Open 405B Instruct
Llama3-OpenBioLLM-70B 0.6661 0.8292 0.8283 Biomedical Open 70B Pretrained
Meta Llama3 70B 0.7379 0.7945 0.8372 General Open 70B Instruct
Phi3 Med. 128K 0.6708 0.8661 0.8266 General Open 14B Instruct
Mixtral 8x7B 0.6033 0.8455 0.8259 General Open 46B Instruct
Gemma2 27B 0.6448 0.8617 0.7874 General Open 27B Instruct
Phi3 Mini 128k 0.5903 0.8804 0.8266 General Open 3.8B Pretrained
Phi3 Mini 4k 0.6036 0.874 0.8186 General Open 3.8B Pretrained
Gemma2 9B 0.6153 0.8435 0.8158 General Open 9B Instruct
Llama3 OpenBioLLM 8B 0.4499 0.8629 0.8246 Biomedical Open 8B Pretrained
Llama3 8B 0.4724 0.8804 0.8344 General Open 8B Pretrained
MetaLlama3.1 8B 0.6189 0.8677 - General Open 8B Instruct
PMCLLAMA 7B 0.4629 0.865 - Biomedical Open 7B Finetuned
JSL MedLlama 8B 0.5726 0.8901 0.8303 Biomedical Open 8B Pretrained
Meditron 7B 0.5102 0.8547 - Biomedical Closed 7B Finetuned
BioMistral 7B 0.4402 0.7938 - Biomedical Open 7B Instruct
Mistral 7B v02 0.4847 0.8709 0.8259 General Open 7.2B Instruct
Mistral 7B v03 0.5084 0.8744 - General Open 7.2B Instruct
Gemma2 2B 0.1728 0.8559 0.817 General Open 2B Instruct
Table 4: Model accuracy on MedQA vs. Expert MCQ, showing performance difference between the two tasks.
Model MedQA AfriMedQA-MCQ-Expert Acc. Difference
Gpt-4o 0.8814 0.7928 -8.86
Gpt-4 0.7989 0.7568 -4.21
Gpt-3.5 Turbo 0.575 0.5629 -1.21
Gpt-4o mini 0.74 0.7176 -2.24
Gemini Ultra 0.7879 0.739 -4.89
Gemini Pro 0.5962 0.6036 0.74
Claude-3.5 Sonnet 0.8327 0.777 -5.57
Claude-3 Opus 0.78 0.7455 -3.45
Claude-3 Sonnet 0.6489 0.6504 0.15
Claude-3 Haiku 0.6709 0.6639 -0.7
Llama3 405B Instruct 0.8068 0.7627 -4.41
Llama3-OpenBioLLM 70B 0.5862 0.6661 7.99
MetaLlama3 70B Instruct 0.7808 0.7379 -4.29
Phi3 Med. 128k 0.6842 0.6708 -1.34
Phi3 Mini 128k 0.575 0.5903 1.53
Phi3 Mini 4k 0.5766 0.6036 2.7
Llama3 8B 0.4973 0.4724 -2.49
MetaLlama3.1 8B Instruct 0.6269 0.6189 -0.8
Llama3 OpenBioLLM 8B 0.4674 0.4499 -1.75
JSL MedLlama 8B 0.6072 0.5726 -3.46
PMC LLAMA 7B 0.509 0.4629 -4.61
Meditron 7B 0.5334 0.5102 -2.32
Mixtral 8x7B 0.6002 0.6033 0.31
Mistral 7B v02 0.5003 0.4847 -1.56
Mistral 7B v03 0.513 0.5084 -0.46
BioMistral 7B 0.4564 0.4402 -1.62
Gemma2 2B 0.3283 0.1728 -15.55
Gemma2 9B 0.6135 0.6153 0.18
Gemma2 27B 0.6209 0.6448 2.39
Table 5: Impact of Explanations: Model accuracy when predicting with explanations vs. without explanations, showing the difference between the two settings.
Model Explanations No Explanations Acc. Difference
Gpt 4 0.8253 0.8247 -0.06
Gpt 4o 0.8276 0.8500 2.24
Gpt 3.5 Turbo 0.6830 0.6890 0.6
Gpt 4o mini - 0.788 78.8
Claude 3.5 Sonnet - 0.8423 -
Claude 3 Sonnet 0.6893 0.733 4.37
Claude 3 Opus 0.7907 0.8110 2.03
Claude 3 Haiku 0.7120 0.7433 3.13
Gemini Pro 0.6310 - -
MedLM 0.7043 - -
Gemini Ultra 0.8003 - -
MedPalm 2 0.7456 - -
MetaLlama3.1 405B - 0.8210 -
OpenBioLLM 70B 0.7280 0.6863 -4.17
MetaLlama3 70B 0.8036 0.8043 0.07
Phi3 Mini 128k 0.6676 0.6813 1.37
Phi3 Med. 128k - 0.7520 75.2
Phi3 Mini 4k 0.6606 0.6803 1.97
OpenBioLLM-8B 0.5193 0.5327 1.34
MetaLlama3 8B 0.6350 0.6003 -3.47
MetaLlama3.1 8B - 0.6933 -
PMCLlama 7B 0.5197 0.5433 2.36
JSLMedLlama3 8B v2.0 0.6606 0.6723 1.17
Meditron 7B 0.5653 0.5807 1.54
BioMistral 7B - 0.5353 -
Mixtral 8x7B v0 0.7203 0.7023 -1.8
Mistral 7B v0.2 0.5510 0.5837 3.27
Table 6: Expert MCQ accuracy by country
Model Kenya Malawi Ghana South Africa Nigeria
Gpt-4o 0.87 0.85 0.82 0.70 0.73
Claude-3.5 Sonnet 0.84 0.86 0.81 0.72 0.71
MetaLlama3 70B 0.79 0.80 0.78 0.63 0.67
Gemini Ultra 0.80 0.78 0.77 0.63 0.67
Llama3 405B Instruct 0.81 0.85 0.79 0.7 0.26
Phi3 Med. 128K 0.73 0.70 0.71 0.59 0.6
MetaLlama3.1 8B 0.69 0.63 0.66 0.61 0.55
Mixtral-8x7B v0.1 0.66 0.67 0.62 0.50 0.55
OpenBioLLM 70B 0.62 0.64 0.63 0.50 0.52
JSL MedLlama3 8B v2.0 0.63 0.57 0.59 0.50 0.54
Meditron 7B 0.54 0.59 0.53 0.39 0.47
BioMistral 7B 0.50 0.40 0.46 0.39 0.41
Avg. Accuracy 0.71 0.70 0.68 0.57 0.48
country num specialties specialties
GH 24 Other (5), Otolaryngology (29), Cardiology (82), Internal Medicine (27), Neurology (116), Infectious Disease (171), Oncology (2), Obstetrics and Gynecology (147), Hematology (15), Plastic Surgery (1), Psychiatry (150), Anesthesiology (1), General Surgery (242), Pulmonary Medicine (62), Pathology (48), Pediatrics (149), Rheumatology (18), Neurosurgery (4), Family Medicine(1), Dermatology (1), Gastroenterology (63), Ophthalmology (35), Nephrology (60), Endocrinology (66)
KE 24 Other (2), Cardiology (28), Internal Medicine (2), Neurology (28), Physical Medicine and Rehabilitation (1), Infectious Disease (12), Oncology (6), Obstetrics and Gynecology (200), Hematology (6), Psychiatry (23), General Surgery (109), Pulmonary Medicine (37), Pediatrics (22), Urology (4), Rheumatology (5), Neurosurgery (1), Family Medicine (1), Geriatrics (1), Dermatology (4), Gastroenterology (15), Orthopedic Surgery (26), Nephrology (18), Endocrinology (10), Emergency Medicine (1), Pathology (0)
MW 27 Other (5), Otolaryngology (11), Cardiology (9), Internal Medicine (41), Neurology (11), Infectious Disease (1), Oncology (28), Radiology (1), Obstetrics and Gynecology (1), Hematology (4), Plastic Surgery (4), Anesthesiology (1), General Surgery (38), Pulmonary Medicine (4), Pathology (2), Pediatrics (74), Urology (4), Rheumatology (2), Neurosurgery (9), Family Medicine (2), Geriatrics (1), Dermatology (3), Gastroenterology (3), Ophthalmology (59), Orthopedic Surgery (6), Nephrology (1), Emergency Medicine (22)
NG 23 Other (20), Otolaryngology (18), Cardiology (53), Internal Medicine (1), Neurology (58), Oncology (2), Radiology (11), Obstetrics and Gynecology (312), Hematology (75), Plastic Surgery (10), Anesthesiology (10), General Surgery (99), Pathology (243), Pediatrics (286), Urology (16), Rheumatology (60), Neurosurgery (17), Family Medicine (1), Dermatology (21), Gastroenterology (49), Ophthalmology (10), Orthopedic Surgery (19), Endocrinology (61)
ZA 1 Pediatrics (54)
Table 7: Counts of Expert MCQ specialties by country
Table 8: Distribution of Number of MCQ Answer Options
num_options counts
5 2718
4 196
2 85
3 1
Country Count Country Count
Nigeria 7577 Tanzania 36
Kenya 2476 Lesotho 33
South Africa 2444 United States 30
Ghana 1572 Zimbabwe 28
Malawi 465 Australia 19
Philippines 320 Botswana 17
Uganda 175 Eswatini 4
Mozambique 77 Zambia 2
Table 9: Counts of Questions Contributed by Country
Table 10: Country Count for Expert Questions
Country Count
Ghana (GH) 1495
Nigeria (NG) 1452
Kenya (KE) 562
Malawi (MW) 347
South Africa (ZA) 54
A.1 Human Evaluation Axes
we collected human evaluations in two categories:
Non-clinicians:
were instructed to provide ratings to LLM responses to determine if answers were relevant (Answer relates to question content), helpful (Answer is helpful and informative), and localized (Answer reflects African local context). These represent the axes of relevance, helpfulness, and bias.
Clinicians:
were instructed to provide ratings to the LLM responses using the following criteria: (a) Correctness: (Correct and consistent with scientific consensus) (b) Harm: (Possibility of harm), (c) Omission: Omission of relevant info, (d)Hallucination: Includes irrelevant, wrong, or extraneous information, (e) African: Requires African local expertise
A.2 Samples of Prompts Used for the Model
We show the output of our post-processing pipeline and how the texts are formatting for several question types. Our prompt design and figure formatting were inspired by (Zheng et al., 2024)
Base MCQ prompt: The following are multiple choice questions (MCQs).
You should directly answer the question by choosing the correct option and then provide a rationale for your answer.
Instruction-tuning MCQ prompt: You are a skilled doctor with years of experience in the medical field in Africa, working in a hospital setting. Your expertise spans a range of conditions from common ailments to complex diseases. As part of your commitment to ongoing medical education, you are evaluating a set of multiple choice questions (MCQs) designed for medical students.
Carefully select the most appropriate answer based on your clinical knowledge.
You should directly answer the question by choosing the correct option and then provide a rationale for your answer.
Base SAQ prompt: The following are short answer questions (SAQs).
You should directly answer the question by providing a short answer and then provide a rationale for your answer.
Instruction-tuning SAQ prompt: You are a skilled doctor with years of experience in the medical field in Africa, working in a hospital setting. Your expertise covers a wide range of conditions from common ailments to complex diseases. You are tasked with evaluating a set of short answer questions (SAQs).
You should provide a concise, direct answer based on your clinical knowledge and experience.
Following your answer, offer a rationale that explains the reasoning behind your response, utilizing medical evidence and current practices to support your explanation.
Base Consumer Queries prompt: The following are open-ended questions.
You should directly answer the question freely.
Instruction-tuning Consumer Queries prompt: You are a skilled doctor with years of experience in the medical field in Africa, working in a hospital setting. Your expertise spans a wide range of conditions from common ailments to complex diseases. You are now addressing a set of open-ended questions designed to explore your medical insights and experiences.
You should answer each question freely, drawing upon your clinical knowledge to provide thorough, informed responses that reflect your understanding of the topics discussed.
Figure 7: Base and Instruction-tuning prompts used for each question.
Question Instruction from Figure 7
Here are some examples:
Question samples and answers as in Figure 2 X number of shots.
###Answer:
Figure 8: Few-shot formatting for in-context learning
###Instruction: The following are multiple choice questions (MCQs). You should directly answer the question by choosing the correct option and then provide a rationale for your answer. {in-context examples from Figure 8 (if few-shots)} ###Question: Which of the following conditions, prevalent in Africa, is caused by infection with the protozoan parasite Trypanosoma brucei and is transmitted to humans through the bite of infected tsetse flies?
###Options:
A. Malaria
B. African trypanosomiasis (sleeping sickness)
C. Chagas disease
D. Leishmaniasis
E. Onchocerciasis
###Answer:
Figure 9: Sample of final Text formatting that is passed into the model
(a) Question Distribution by Country
(b) Distribution by Gender
(c) Distribution by Question Type
Figure 10: Dataset Distributions
Figure 11: MCQ User Interface
Figure 12: MCQ Review Interface with contributor and reviewer name redacted
Figure 13: Dataset Annotation Tool
Figure 14: MCQ/SAQ instructions
Table 11: Specialty Count for All and Expert Questions
Specialty All Expert
Obs. and Gynecology 824 660
Pediatrics 747 585
General Surgery 757 488
Pathology 381 293
Neurology 310 213
Infectious Disease 539 184
Psychiatry 299 173
Cardiology 258 172
Endocrinology 236 137
Gastroenterology 225 130
Allergy and Immunology 217 -
Ophthalmology 202 104
Pulmonary Medicine 231 103
Hematology 211 100
Rheumatology 171 85
Nephrology 163 79
Internal Medicine 238 71
Otolaryngology 158 58
Orthopedic Surgery 161 51
Oncology 135 38
Other 125 32
Family Medicine 289 -
Neurosurgery 107 31
Radiology 107 -
Phy. Med. and Rehab. 101 -
Dermatology 126 29
Urology 134 24
Emergency Medicine 123 23
Plastic Surgery 132 15
Anesthesiology 115 -
Geriatrics 91 -
Medical Genetics 86 -
Table 12: MCQs Post-Processing Errors Count for base prompts
Model Errors Count
Llama3-OpenBioLLM-8B 435
Claude-3-sonnet-20240229 273
Claude-3-haiku-20240307 164
Llama3-OpenBioLLM-70B 163
Claude-3-opus-20240229 162
GPT-3.5-turbo 120
Phi3-mini-128 91
GPT-4o 80
MedPALM2 40
LLama-3-8b 26
GPT-4-turbo 17
GPT-4 10
Meta-Llama-3-70B-Instruct 6
Mixtral-8x7B-Instruct-v0.1 5
Gemini ultra 4
Meta-Llama-3-8B-Instruct 2
Gemini Pro 1
MedLM 1
Table 13: MCQs Post-Processing Errors Count for Instruct Prompts
Model Name Post-Processing Errors Count
Llama3-OpenBioLLM-8B 456
Meta-Llama-3-8B-Instruct 191
Mistral-7b 167
Claude-3-opus-20240229 162
Llama3-OpenBioLLM-70B 153
MedPALM2 60
Phi3-mini-4 43
Mixtral-8x7B-Instruct-v0.1 18
GPT-4o 11
Meta-Llama-3-70B-Instruct 8
MedLM 1
Gemini ultra 1
Table 14: HyperParameters Used for all models
Model Names Temperature Batch Size Max (New) Tokens
Claude 3 Haiku 0.2 1 1000
Claude 3 Opus 0.2 1 1000
Claude 3 sonnet 0.2 1 1000
Gemini Pro 0.9 32 -
Gemini Ultra 0.7 16 -
GPT 3.5 turbo - 1 -
GPT-4 - 1 -
GPT-4 turbo - 1 -
GPT-40 - 1 -
JSL-MedLlama-3-8B-v2.0 0.7 - 256
Llama-3-70B-Instruct - - -
Llama3 8B - 1 300
Llama-3-8B-Instruct - - -
MedLM 0.2 16 -
MedPalm 2 - - -
Mistral-7B-Instruct-v0.2 - 1 -
Mixtral-8x7B-Instruct-v0.1 - 1 500
OpenBioLLM-70B-Instruct - - -
OpenBioLLM-8B-Instruct - - -
Phi-3-mini-128k-instruct-3.8B 0.0 1 500
Phi-3-mini-4k-instruct-3.8B 0.0 1 500
  • "-" indicates default hyperparameters.
Table 15: SAQ Human Eval: Clinician blind evaluations of human and LLM answers showing counts and mean ratings.
model count Correct Harm Hallucinations Omission
claude-3-sonnet 176 4.676 1.023 1.023 1.142
gpt-3.5-turbo 182 4.692 1.005 1.044 1.192
gpt-4 195 4.605 1.026 1.082 1.169
gpt-4-turbo 187 4.668 1.064 1.102 1.166
gpt-4o 178 4.635 1.034 1.045 1.185
human 197 4.442 1.051 1.061 1.533
jsl-med-llama-8b 182 4.588 1.038 1.055 1.308
llama-3-8b 176 4.273 1.080 1.233 1.557
mistral-7b 193 4.751 1.016 1.104 1.052
phi3-mini-128 158 4.608 1.000 1.032 1.177
phi3-mini-4 175 4.714 1.023 1.017 1.109
Table 16: Clinician blind evaluations of human and LLM answers showing counts and mean ratings for the Correct and consistent with scientific consensus axis for all rated MCQs and the percentage change on questions human evaluators rated greater than 1 for Requires African local expertise.
model all cnt Afr cnt all mean Afr mean $\Delta$%
Llama3-OpenBioLLM-8B 133 5 4.211 3.600 -14.500
Llama3-OpenBioLLM-70B 169 7 4.444 3.857 -13.201
medpalm2 56 3 4.607 4.000 -13.178
gemini-ultra 36 5 4.583 4.200 -8.364
claude-3-haiku-20240307 83 3 4.639 4.333 -6.580
Mixtral-8x7B-Instruct-v0.1 235 9 4.434 4.222 -4.777
Meta-Llama-3-8B-Instruct 166 11 4.452 4.273 -4.023
claude-3-opus-20240229 79 4 4.658 4.500 -3.397
claude-3-sonnet-20240229 857 158 4.693 4.595 -2.092
human 1430 314 4.631 4.586 -0.967
gpt-4-turbo 1232 189 4.695 4.693 -0.036
phi3-mini-128 1127 171 4.650 4.649 -0.008
gpt-4 1248 211 4.599 4.602 0.073
mistral-7b 773 156 4.604 4.641 0.801
gpt-4o 1884 370 4.720 4.762 0.899
gpt-3.5-turbo 1202 166 4.656 4.699 0.928
phi3-mini-4 1159 177 4.651 4.723 1.561
Meta-Llama-3-70B-Instruct 150 8 4.520 4.625 2.323
llama-3-8b 1220 174 4.284 4.402 2.751
jsl-med-llama-8b 1197 183 4.420 4.568 3.350
medlm 29 2 4.414 5.000 13.281
Table 17: Consumer blind evaluations of LLM answers showing counts and mean ratings.
model count Relevant Helpful+Informative Localized
claude-3-haiku 1430 4.448 4.349 1.466
claude-3-opus 1407 4.465 4.387 1.425
claude-3-sonnet 2755 4.531 4.446 1.643
gemini-pro 1823 4.440 4.332 1.231
gemini-ultra 967 4.074 3.815 1.532
gpt-3.5-turbo 1429 4.319 3.946 1.824
gpt-4 1347 4.324 4.143 1.869
gpt-4-turbo 1658 4.504 4.257 1.800
gpt-4o 1338 4.340 4.243 1.938
jsl-med-llama-8b 1337 4.196 4.016 1.818
llama-3-8b 1374 4.097 3.832 1.845
medlm 998 4.200 4.060 1.441
medpalm2 959 3.938 3.614 1.501
mistral-7b 1343 4.331 4.207 1.842
phi3-mini-128 1361 4.212 4.139 1.882
phi3-mini-4 1343 4.416 4.343 1.826
Table 18: Clinician blind evaluations of human and LLM answers showing number of times Possibility of harm was rated greater than 1 for all rated MCQs and the percentage change on MCQs human evaluators rated greater than 1 for Requires African local expertise. Due to the random order of evaluations and order of model result submissions, models received an unequal number of ratings at the project deadline. How to read this table: for example, of the 1,248 human ratings for gpt-4, 28 responses (2.2%) were rated as having the possibility of harm. Of the 1,248 ratings, 211 were also rated as requiring African local expertise. Of these 211, 7 MCQ answers (1.9%) had the possibility of harm, a 15.5% improvement
model cnt all all pct cnt Afr all Afr Afr pct $\Delta$%
medpalm2 1 56 1.786 1 3 33.333 1766.349
human 39 1430 2.727 17 314 5.414 98.533
llama-3-8b 52 1220 4.262 7 174 4.023 -5.608
mistral-7b 13 773 1.682 5 156 3.205 90.547
claude-3-sonnet 16 857 1.867 5 158 3.165 69.523
gpt-4-turbo 18 1232 1.461 5 189 2.646 81.109
phi3-mini-4 19 1159 1.639 4 177 2.260 37.889
jsl-med-llama-8b 35 1197 2.924 4 183 2.186 -25.239
gpt-4 28 1248 2.244 4 211 1.896 -15.508
phi3-mini-128 18 1127 1.597 2 171 1.170 -26.738
gpt-4o 23 1884 1.221 3 370 0.811 -33.579
gpt-3.5-turbo 16 1202 1.331 1 166 0.602 -54.771
Meta-Llama-3-8B-Instruct 1 166 0.602 0 11 $0$
Llama3-OpenBioLLM-8B 2 133 1.504 0 5 $0$
Mixtral-8x7B-Instruct-v0.1 1 235 0.426 0 9 $0$
claude-3-haiku 0 83 0.000 0 3 $0$
claude-3-opus 0 79 0.000 0 4 $0$
gemini-ultra 0 36 0.000 0 5 $0$
Llama3-OpenBioLLM-70B 3 169 1.775 0 7 $0$
gemini-pro 0 40 0.000 0 0 $0$
Meta-Llama-3-70B-Instruct 0 150 0.000 0 8 $0$
medlm 1 29 3.448 0 2 $0$
Table 19: Clinician blind evaluations of human and LLM answers showing hallucinations, i.e. number of times Includes irrelevant, wrong, or extraneous information was rated greater than 1 for all rated MCQs and the percentage change on MCQs human evaluators rated greater than 1 for Requires African local expertise.
model cnt all all % cnt Afr all Afr Afr % $\Delta$%
llama-3-8b 117 1220 9.590 11 174 6.322 -34.077
jsl-med-llama-8b 97 1197 8.104 9 183 4.918 -39.314
mistral-7b 47 773 6.080 9 156 5.769 -5.115
Meta-Llama-3-8B-Instruct 10 166 6.024 1 11 9.091 50.913
Llama3-OpenBioLLM-8B 7 133 5.263 1 5 20.000 280.011
human 72 1430 5.035 16 314 5.096 1.212
gpt-4o 89 1884 4.724 18 370 4.865 2.985
Meta-Llama-3-70B-Instruct 7 150 4.667 1 8 12.500 167.838
claude-3-sonnet 36 857 4.201 9 158 5.696 35.587
phi3-mini-4 43 1159 3.710 5 177 2.825 -23.854
gpt-4 45 1248 3.606 7 211 3.318 -7.987
gpt-4-turbo 44 1232 3.571 5 189 2.646 -25.903
Llama3-OpenBioLLM-70B 6 169 3.550 0 7 $0$
gpt-3.5-turbo 36 1202 2.995 5 166 3.012 0.568
gemini-pro 1 40 2.500 0 0 $0$
phi3-mini-128 26 1127 2.307 5 171 2.924 26.745
claude-3-opus 1 79 1.266 0 4 $0$
Mixtral-8x7B-Instruct-v0.1 2 235 0.851 0 9 $0$
claude-3-haiku 0 83 0.000 0 3 $0$
gemini-ultra 0 36 0.000 0 5 $0$
medpalm2 0 56 0.000 0 3 $0$
medlm 0 29 0.000 0 2 $0$
Table 20: Clinician blind evaluations of human and LLM answers showing ommissions, i.e. number of times Omission of relevant info was rated greater than 1 for all rated MCQs and the percentage change on MCQs human evaluators rated greater than 1 for Requires African local expertise.
model cnt all all % cnt Afr all Afr Afr % $\Delta$%
Llama3-OpenBioLLM-8B 36 133 27.068 1 5 20.000 -26.112
llama-3-8b 264 1220 21.639 39 174 22.414 3.581
jsl-med-llama-8b 201 1197 16.792 29 183 15.847 -5.628
Llama3-OpenBioLLM-70B 24 169 14.201 2 7 28.571 101.190
human 201 1430 14.056 56 314 17.834 26.878
gpt-4 146 1248 11.699 18 211 8.531 -27.079
gpt-4-turbo 120 1232 9.740 19 189 10.053 3.214
phi3-mini-128 107 1127 9.494 16 171 9.357 -1.443
mistral-7b 73 773 9.444 15 156 9.615 1.811
gpt-3.5-turbo 113 1202 9.401 17 166 10.241 8.935
phi3-mini-4 108 1159 9.318 20 177 11.299 21.260
claude-3-sonnet 72 857 8.401 21 158 13.291 58.207
gpt-4o 139 1884 7.378 25 370 6.757 -8.417
medpalm2 4 56 7.143 1 3 33.333 366.653
gemini-pro 2 40 5.000 0 0 $0$
Meta-Llama-3-70B-Instruct 4 150 2.667 0 8 $0$
Meta-Llama-3-8B-Instruct 3 166 1.807 0 11 $0$
claude-3-opus 1 79 1.266 0 4 $0$
Mixtral-8x7B-Instruct-v0.1 2 235 0.851 1 9 11.111 1205.640
claude-3-haiku 0 83 0.000 0 3 $0$
gemini-ultra 0 36 0.000 0 5 $0$
medlm 0 29 0.000 0 2 $0$
Table 21: Specialty Human Eval: Clinician ratings for all model responses across specialties showing counts of each specialty, mean of correctness rating, and MCQ counts with percentages where model responses demonstrated the possibility of harm, hallucinations, and omission of relevant information, i.e. ratings greater than 1 for relevant criteria. Bold represents bottom 2 specialties with worst performance across all models for each evaluation axis
specialty count correct # harm (%) # hallucination (%) # omission (%)
Obstetrics and Gynecology 482 4.620 10 (2.07) 29 (6.02) 43 (8.92)
Internal Medicine 323 4.635 9 (2.79) 19 (5.88) 24 (7.43)
Infectious Disease 319 4.680 3 (0.94) 5 (1.57) 9 (2.82)
Cardiology 280 4.736 4 (1.43) 11 (3.93) 8 (2.86)
Pediatrics 274 4.661 2 (0.73) 10 (3.65) 18 (6.57)
Endocrinology 254 4.701 3 (1.18) 13 (5.12) 8 (3.15)
Neurology 217 4.765 3 (1.38) 4 (1.84) 8 (3.69)
Other 215 4.521 3 (1.40) 6 (2.79) 11 (5.12)
Gastroenterology 177 4.616 6 (3.39) 13 (7.34) 17 (9.60)
Hematology 154 4.766 2 (1.30) 1 (0.65) 7 (4.55)
General Surgery 144 4.597 8 (5.56) 12 (8.33) 13 (9.03)
Pathology 108 4.731 2 (1.85) 1 (0.93) 9 (8.33)
Family Medicine 107 4.748 0 (0.00) 5 (4.67) 6 (5.61)
Orthopedic Surgery 97 4.680 0 (0.00) 7 (7.22) 6 (6.19)
Pulmonary Medicine 94 4.755 3 (3.19) 4 (4.26) 10 (10.64)
Ophthalmology 88 4.602 2 (2.27) 3 (3.41) 10 (11.36)
Oncology 87 4.678 2 (2.30) 4 (4.60) 4 (4.60)
Nephrology 79 4.570 1 (1.27) 6 (7.59) 8 (10.13)
Emergency Medicine 71 4.662 0 (0.00) 4 (5.63) 4 (5.63)
Anesthesiology 69 4.725 3 (4.35) 2 (2.90) 5 (7.25)
Urology 56 4.714 0 (0.00) 2 (3.57) 4 (7.14)
Psychiatry 54 4.667 1 (1.85) 3 (5.56) 3 (5.56)
Dermatology 50 4.700 1 (2.00) 1 (2.00) 2 (4.00)
Neurosurgery 48 4.438 1 (2.08) 2 (4.17) 5 (10.42)
Allergy and Immunology 43 4.605 1 (2.33) 2 (4.65) 3 (6.98)
Otolaryngology 35 4.743 1 (2.86) 0 (0.00) 2 (5.71)
Radiology 23 4.478 0 (0.00) 2 (8.70) 1 (4.35)
Medical Genetics 18 5.000 0 (0.00) 0 (0.00) 0 (0.00)
Plastic Surgery 17 4.706 0 (0.00) 1 (5.88) 1 (5.88)
Rheumatology 16 4.625 0 (0.00) 1 (6.25) 1 (6.25)
Geriatrics 11 5.000 0 (0.00) 0 (0.00) 0 (0.00)

### v3

@@
 However, their effectiveness in the Global South, especially across the African continent, remains to be established.
-In this work, we introduce AfriMed-QA 22footnotetext: Our dataset is available at:
-https://huggingface.co/datasets/intronhealth/afrimedqa_v2
-, the first large-scale Pan-African English multi-specialty medical Question-Answering (QA) dataset, 15,000 questions (open and closed-ended) sourced from over 60 medical schools across 16 countries, covering 32 medical specialties.
+In this work, we introduce AfriMed-QA, the first large-scale Pan-African English multi-specialty medical Question-Answering (QA) dataset, 15,000 questions (open and closed-ended) sourced from over 60 medical schools across 16 countries, covering 32 medical specialties.
 We further evaluate 30 LLMs across multiple axes including correctness and demographic bias.
@@
 (2019, 2021), but these may not translate to out-of-distribution datasets, such as those from African countries.
-To address this gap, we introduce AfriMed-QA, a dataset of 15,275 English, clinically diverse questions and answers, 4,000+ expert multiple-choice questions (MCQs) with answers, over 1,200 open-ended short answer (SAQs) with long-form answers, and 10,000 consumer queries (CQ), to rigorously assess LLM performance for correctness and demographic bias.
+To address this gap, we introduce AfriMed-QA, 22footnotetext: Available at:
+https://huggingface.co/datasets/intronhealth/afrimedqa_v2
+a dataset of 15,275 English, clinically diverse questions and answers, 4,000+ expert multiple-choice questions (MCQs) with answers, over 1,200 open-ended short answer (SAQs) with long-form answers, and 10,000 consumer queries (CQ), to rigorously assess LLM performance for correctness and demographic bias.
 We evaluate 30 large, small, open, closed, biomedical, and general LLMs using quantitative and qualitative approaches.
 While the development of the dataset is still in progress, this work establishes a foundation for acquiring diverse and representative health benchmark datasets across LMICs.
+The dataset is released under a CC-BY-NC-SA 4.0 license.
 2 Related Work
@@
 These studies underscore the importance of diverse, high-quality datasets and human assessments to capture the nuances of real-world medical applications.
-Reddy et al.
 Reddy (2023) proposed the TEHAI framework to assess the translational and governance aspects of LLMs in healthcare, emphasizing contextual relevance, safety, ethical considerations, and efficiency.
 Our evaluation extends TEHAI and work from Singhal et al.
-Singhal et al.
 (2022b), and Pfohl et al.
-Pfohl et al.
 (2024) by incorporating specialty and region-specific dimensions, ensuring a comprehensive assessment of LLMs.
@@
 (2020).
-Figure 3: AfriMedQA: Expert-MCQ accuracy, SAQ, and CQ Bertscore
+Figure 3: AfriMedQA: Expert-MCQ accuracy, SAQ, and CQ Bertscore, sorted by Expert-MCQ Accuracy
 (a) Top: MedQA vs AfriMedQA MCQ Accuracy. Bottom: Effect of Explanations
@@
 We recommend its use for benchmarking and finetuning, recognizing its potential to drive advancements in medical LLMs that are culturally attuned to the unique needs of African populations and other regions in the Global South.
-Acknowledgements
-Support for this project was provided by Google Research, PATH, and the Bill & Melinda Gates Foundation [Gates INV-068056].
-We appreciate the excellent execution by the Intron Health Data Team who managed this pan-African data collection effort including recruiting contributors, developing custom UIs, quality reviews, and multi-currency contributor payments.
-We appreciate the invaluable support provided by the BioRAMP researchers, whose collaboration and insights have been fundamental to our research.
-We also extend our thanks to the Masakhane Community for their expert contributions, the Sisonke Biotik community for their enthusiastic participation and commitment, and the Federation of African Medical Students’ Associations [FAMSA] for their relentless recruitment and community engagement.
-These partnerships have been instrumental in enhancing the scope and impact of our work, and we appreciate their dedication to advancing knowledge in our field.
 Appendix A Appendix

### v4

@@
 AfriMed-QA: A Pan-African, Multi-Specialty, Medical Question-Answering Benchmark Dataset
-Tobi Olatunji1,2,4, Charles Nimo2, Abraham Owodunni1,3, Tassallah Abdullahi4,12, Emmanuel Ayodele1, Mardhiyah Sanni1,4, Chinemelu Aka1, Folafunmi Omofoye4, Foutse Yuehgoh4, Timothy Faniran4, Bonaventure F.
+Tobi Olatunji1,2,4 thanks: These authors contributed equally to this work., Charles Nimo2  [^1] Available at:
+https://huggingface.co/datasets/intronhealth/afrimedqa_v2
+, Abraham Owodunni1,3  [^1] Available at:
+https://huggingface.co/datasets/intronhealth/afrimedqa_v2
+Tassallah Abdullahi4,12, Emmanuel Ayodele1, Mardhiyah Sanni1,4, Chinemelu Aka1, Folafunmi Omofoye4, Foutse Yuehgoh4, Timothy Faniran4, Bonaventure F.
 P.
-Dossou4,10, Moshood Yekini4, Jonas Kemp6, Katherine Heller6, Jude Chidubem Omeke4, Chidi Asuzu MD4, Naome A.
-Etori4,11, Aimérou Ndiaye5, Ifeoma Okoh5, Evans Doe Ocansey5, Wendy Kinara7, Michael Best2, Irfan Essa2,6, Stephen Edward Moore8, Chris Fourie9, Mercy Nyamewaa Asiedu6♠ 1Intron, 2Georgia Institute of Technology, 3The Ohio State University, 4BioRAMP, 5Masakhane, 6Google Research, 7Kenyatta University, 8University of Cape Coast, 9SisonkeBiotik, 10MILA Quebec, 11University of Minnesota, 12Brown University tobi@intron.io
-11footnotetext: Authors marked with $\spadesuit$ are senior authors.
+Dossou4,10,13, Moshood Yekini4, Jonas Kemp6, Katherine Heller6, Jude Chidubem Omeke4, Chidi Asuzu MD4, Naome A.
+Etori4,11, Aimérou Ndiaye5, Ifeoma Okoh5, Evans Doe Ocansey5, Wendy Kinara7, Michael Best2, Irfan Essa2,6, Stephen Edward Moore8, Chris Fourie9, Mercy Nyamewaa Asiedu6 thanks: Senior authors.
+1Intron, 2Georgia Institute of Technology, 3The Ohio State University, 4BioRAMP, 5Masakhane, 6Google Research, 7Kenyatta University, 8University of Cape Coast, 9SisonkeBiotik, 10MILA Quebec, 11University of Minnesota, 12Brown University 13McGill University tobi@intron.io
 1 Introduction
@@
 Table 1: Comparative analysis of the AfriMed-QA dataset with health datasets.
-Feature AfriMed-QA BioASQ Krithara et al.
-(2023) MedQA-SWE Hertzberg and Lokrantz (2024) MedQA Jin et al.
-(2020) MedMCQA Pal et al.
-(2022b) PubMedQA Jin et al.
-(2019) MMLU Hendrycks et al.
-(2020) HealthSearchQA Singhal et al.
-(2022a)
+Feature AfriMed-QA
+BioASQ
+Krithara et al. (2023)
+MedQA-SWE
+Hertzberg and Lokrantz (2024)
+MedQA
+Jin et al. (2020)
+MedMCQA
+Pal et al. (2022b)
+PubMedQA
+Jin et al. (2019)
+MMLU
+Hendrycks et al. (2020)
+HealthSearchQA
+Singhal et al. (2022a)
 Dataset Size 15,275 4,721 3,180 12,723 193,155 1,000 15,908 3,173
@@
 (2019, 2021), but these may not translate to out-of-distribution datasets, such as those from African countries.
-To address this gap, we introduce AfriMed-QA, 22footnotetext: Available at:
-https://huggingface.co/datasets/intronhealth/afrimedqa_v2
-a dataset of 15,275 English, clinically diverse questions and answers, 4,000+ expert multiple-choice questions (MCQs) with answers, over 1,200 open-ended short answer (SAQs) with long-form answers, and 10,000 consumer queries (CQ), to rigorously assess LLM performance for correctness and demographic bias.
+To address this gap, we introduce AfriMed-QA, a dataset of 15,275 English, clinically diverse questions and answers, 4,000+ expert multiple-choice questions (MCQs) with answers, over 1,200 open-ended short answer (SAQs) with long-form answers, and 10,000 consumer queries (CQ), to rigorously assess LLM performance for correctness and demographic bias.
 We evaluate 30 large, small, open, closed, biomedical, and general LLMs using quantitative and qualitative approaches.
@@
 3.1 Data Collection
-We adapted a web-based platform previously developed by Intron Health  [^1] Intron Health’s biomedical crowd-sourcing platform https://speech.intron.health to crowd-source accented and multilingual clinical speech data at scale across Africa Olatunji et al.
+We adapted a web-based platform previously developed by Intron Health 22footnotetext: Intron Health’s biomedical crowd-sourcing platform https://speech.intron.health to crowd-source accented and multilingual clinical speech data at scale across Africa Olatunji et al.
 (2023).
@@
 Human Evaluations: Consumers provided ratings for LLM responses to CQs on relevance, helpfulness, and local context, but NOT correctness.
-Due of the higher expertise required, only confirmed clinicians were granted access to projects rating MCQ, SAQ, and CQ answers.
+Due to the higher expertise required, only confirmed clinicians were granted access to projects rating MCQ, SAQ, and CQ answers.
 Clinician status was confirmed through submitted credentials and background checks reviewing publicly available information about them.
@@
 4.1 LLM Selection
-We evaluate 30 LLMs (Table 3) including open-source and proprietary LLMs, general-purpose and biomedical LLMs, mixture-of-experts, and models of varying sizes from 3B to over 540B parameters Chowdhery et al.
-(2023)).
+We evaluate 30 LLMs (Table 3) including open-source and proprietary LLMs, general-purpose and biomedical LLMs, mixture-of-experts, and models of varying sizes from 3B to over 540B parameters.
 Of these, 13 are proprietary while 17 are open-source.
@@
 For multiple-choice tasks, accuracy is measured by comparing LLM’s single-letter answer choice [A,B,C,D,E] with the reference.
-For open-ended questions, semantic similarity is measured using BertScore, comparing the generated response from the language model against a reference answer Zhang* et al.
-(2020).
+For open-ended questions, we evaluate semantic similarity using BERTScore Zhang* et al.
+(2020) and QuestEval Scialom et al.
+(2021), both comparing the generated response from the language model against a reference answer, and sentence-level structural overlap using ROUGE-Lsum Lin (2004), which likewise compares the generated response against its reference.
 Figure 3: AfriMedQA: Expert-MCQ accuracy, SAQ, and CQ Bertscore, sorted by Expert-MCQ Accuracy
@@
 Ratings were on a 5-point scale representing the extent to which the criteria were met.
-1 represents “No" or “completely absent", and 5 represents “Yes" or “Absolutely present".
+1 represents "No" or "completely absent", and 5 represents "Yes" or "Absolutely present".
 Raters were blinded to the answer source (model name or human).
@@
 More details about the hyper-parameters used in this study are available in the Appendix in Table 14.
-Aligning with our community participatory design approach, experiments by multiple collaborators ran on various hardware types including NVIDIA L4, NVIDIA T4, and A100.
+Aligning with our community participatory design approach, experiments by multiple collaborators ran on various hardware types including 1 of either NVIDIA L4, NVIDIA T4, or A100.
 5 Results
@@
 This finding is counter-intuitive and requires further investigation.
+6.3 Understanding the domain shift
+Although it seems intuitive that medical knowledge is universal, our findings highlight key reasons for regional differences in LLM Performance.
+Questions require not only general biomedical knowledge but also contextual understanding of African-specific disease patterns, local health challenges, and socio-cultural factors.
+For example, children in specific communities may be required to take vaccines (e.g.
+HPV) earlier or later based on logistical access challenges.
+Diseases may also involve region-specific appearance of symptoms (e.g.
+appropriate color for skin lesion) and clinical presentation (e.g.
+cancer patients typically seek professional medical care much later than Western countries, affecting how they are managed on the first visit), and management may reflect available medications, treatment options, and diagnostic equipment.
+Although there is significant medical knowledge overlap globally, regional variations exist, necessitating physician board certification in each country.
+Such local variations make questions more challenging for LLMs trained mainly on Western medical data.
 6.4 Evidence of progress in LLM reasoning abilities
@@
 Country-Specialty counts are detailed in Appendix Tables 7, 6, and 10.
+6.8 Limitations of automated metrics for evaluating SAQ and CQ answers
+As shown by the narrow range of BERTScore values in Appendix Table 3 (all LLMs cluster between 0.86 and 0.89), its utility in this context is limited—a problem also identified in the literature Hanna and Bojar (2021); Celikyilmaz et al.
+(2020); Zhang et al.
+(2024).
+Appendix Table 3 also reports ROUGE-Lsum and QuestEval scores.
+ROUGE-Lsum spans a much broader interval (0.009–0.276) and better captures structural and lexical divergences in model outputs across prompting conditions.
+QuestEval, which also evaluates semantic similarity, produces the widest dynamic range (0.19–0.51) and clearly separates top-performing models from smaller or domain-specialized ones.
+This variability demonstrates that no single automated metric reliably captures both the correctness and completeness of free-text medical responses Tu et al.
+(2024), even when metrics are combined.
+We therefore defer analysis of open-ended answers to future human evaluations where clinicians can better discriminate between semantically similar and clinically correct answers.
 6.9 Positive and negative insights from LLM explanations
@@
 We quantitatively and qualitatively evaluate thirty open and closed-source LLMs demonstrating performance variability across specialties, and geographies.
+8 Ethical Considerations
+The public release of healthcare-related datasets often raises important ethical concerns, especially regarding privacy and consent.
+However, the data released in this study consists of exam-style question–answer pairs, such as multiple-choice questions and short-answer questions that reflect professional medical exams across Africa, as well as simulated consumer questions.
+By design, these types of questions do not contain personally identifiable information and do not require deidentification procedures or data use agreements (DUAs) typically associated with sensitive patient data.
+Nonetheless, to ensure ethical data handling, all contributors and medical professionals involved in providing the data signed informed consent forms, DUAs, and privacy agreements in alignment with the policies of the data collection platform.
 9 Limitations and Future Work
@@
 Appendix A Appendix
-Table 3: Zero-Shot Expert Evaluation of Large Language Models: Accuracy and BERT Scores on Multiple Choice Questions, Short Answer Questions, and Consumer Queries.
+Table 3: Zero-Shot Expert Evaluation of Large Language Models: Accuracy, BERT Scores, ROUGE-Lsum (R-L) Scores, and QuestEval (QE) scores, on Multiple Choice Questions, Short Answer Questions, and Consumer Queries.
 Some model names are shortened for brevity.
-Model MCQ-Expert SAQ Bertscore CQ Bertscore Domain Access Size Training
-Gpt-4 0.7568 0.8727 0.8385 General Closed 1.8T Instruct
-Gpt-4o 0.7928 0.8825 0.8254 General Closed 12B Instruct
-Gpt-3.5 Turbo 0.5629 0.8813 0.0963 General Closed 175B Instruct
-Gpt-4o mini 0.7176 0.8808 - General Closed $\sim$8B Instruct
-Claude-3.5 Sonnet 0.777 0.8574 - General Closed - Instruct
-Claude-3 Sonnet 0.6504 0.8719 0.8141 General Closed 70B Instruct
-Claude-3 Opus 0.7455 0.8696 0.8172 General Closed 2T Instruct
-Claude 3 Haiku 0.6639 0.8656 0.8163 General Closed 2T Instruct
-Gemini Pro 0.6036 0.8601 0.8213 General Closed 540B Instruct
-MedLM 0.6036 0.8633 0.8303 Biomedical Closed - Instruct
-Gemini Ultra 0.739 0.8716 0.8362 General Closed 1.56T Instruct
-MedPalm2 - 0.8716 0.8379 Biomedical Closed 540B Instruct
-llama3-405b-instruct-maas 0.7627 - General Open 405B Instruct
-Llama3-OpenBioLLM-70B 0.6661 0.8292 0.8283 Biomedical Open 70B Pretrained
-Meta Llama3 70B 0.7379 0.7945 0.8372 General Open 70B Instruct
-Phi3 Med. 128K 0.6708 0.8661 0.8266 General Open 14B Instruct
-Mixtral 8x7B 0.6033 0.8455 0.8259 General Open 46B Instruct
-Gemma2 27B 0.6448 0.8617 0.7874 General Open 27B Instruct
-Phi3 Mini 128k 0.5903 0.8804 0.8266 General Open 3.8B Pretrained
-Phi3 Mini 4k 0.6036 0.874 0.8186 General Open 3.8B Pretrained
-Gemma2 9B 0.6153 0.8435 0.8158 General Open 9B Instruct
-Llama3 OpenBioLLM 8B 0.4499 0.8629 0.8246 Biomedical Open 8B Pretrained
-Llama3 8B 0.4724 0.8804 0.8344 General Open 8B Pretrained
-MetaLlama3.1 8B 0.6189 0.8677 - General Open 8B Instruct
-PMCLLAMA 7B 0.4629 0.865 - Biomedical Open 7B Finetuned
-JSL MedLlama 8B 0.5726 0.8901 0.8303 Biomedical Open 8B Pretrained
-Meditron 7B 0.5102 0.8547 - Biomedical Closed 7B Finetuned
-BioMistral 7B 0.4402 0.7938 - Biomedical Open 7B Instruct
-Mistral 7B v02 0.4847 0.8709 0.8259 General Open 7.2B Instruct
-Mistral 7B v03 0.5084 0.8744 - General Open 7.2B Instruct
-Gemma2 2B 0.1728 0.8559 0.817 General Open 2B Instruct
+Additionally, Gen.
+and Biomed.
+here is short for General and Biomedical respectively.
+Model MCQ Expert SAQ BERT CQ BERT SAQ R-L CQ R-L SAQ QE CQ QE Domain Access Size Type
+Gpt-4 0.7568 0.8727 0.8385 0.2054 0.0018 0.4868 0.2352 Gen. Closed 1.8T Instruct
+Gpt-4o 0.7928 0.8825 0.8254 0.2519 0.0014 0.4959 0.2352 Gen. Closed 12B Instruct
+Gpt-3.5 Turbo 0.5629 0.8813 0.0963 0.2452 0.0015 0.5072 0.2353 Gen. Closed 175B Instruct
+Gpt-4o mini 0.7176 0.8808 - 0.2311 - 0.4946 - Gen. Closed $\sim$8B Instruct
+Claude-3.5 Sonnet 0.777 0.8574 - - - 0.5071 - Gen. Closed - Instruct
+Claude-3 Sonnet 0.6504 0.8719 0.8141 0.1984 0.0010 0.4976 0.2352 Gen. Closed 70B Instruct
+Claude-3 Opus 0.7455 0.8696 0.8172 0.1892 0.0010 0.4904 0.2354 Gen. Closed 2T Instruct
+Claude 3 Haiku 0.6639 0.8656 0.8163 0.1929 0.0010 0.5004 0.2352 Gen. Closed 2T Instruct
+Gemini Pro 0.6036 0.8601 0.8213 0.2061 0.0012 0.4370 0.2347 Gen. Closed 540B Instruct
+MedLM 0.6036 0.8633 0.8303 0.1991 0.0014 0.4443 0.2380 Biomed. Closed - Instruct
+Gemini Ultra 0.739 0.8716 0.8362 0.2617 0.0018 0.4368 0.2347 Gen. Closed 1.56T Instruct
+MedPalm2 - 0.8716 0.8379 0.2253 0.0018 0.4304 0.2352 Biomed. Closed 540B Instruct
+llama3-405B 0.7627 - 0.1096 - 0.3744 - Gen. Open 405B Instruct
+OpenBioLLM 70B 0.6661 0.8292 0.8283 0.1866 - 0.3107 - Biomed. Open 70B Instruct
+Meta Llama3 70B 0.7379 0.7945 0.8372 0.0089 - 0.2331 - Gen. Open 70B Instruct
+Phi3 Med. 128K 0.6708 0.8661 0.8266 0.2432 - 0.4999 - Gen. Open 14B Instruct
+Mixtral 8x7B 0.6033 0.8455 0.8259 0.2045 - 0.3783 - Gen. Open 46B Instruct
+Gemma2 27B 0.6448 0.8617 0.7874 - - - - Gen. Open 27B Instruct
+Phi3 Mini 128k 0.5903 0.8804 0.8266 0.2421 0.0014 0.4868 0.2351 Gen. Open 3.8B Pretrained
+Phi3 Mini 4k 0.6036 0.874 0.8186 0.2098 0.0011 0.4894 0.2352 Gen. Open 3.8B Pretrained
+Gemma2 9B 0.6153 0.8435 0.8158 - - - - Gen. Open 9B Instruct
+OpenBioLLM 8B 0.4499 0.8629 0.8246 0.2017 - 0.3877 - Biomed. Open 8B Instruct
+Llama3 8B 0.4724 0.8804 0.8344 0.2421 0.0021 0.4868 0.2336 Gen. Open 8B Pretrained
+MetaLlama3.1 8B 0.6189 0.8677 - 0.1901 - 0.4817 - Gen. Open 8B Instruct
+PMC-Llama 7B 0.4629 0.865 - 0.2194 - 0.3853 - Biomed. Open 7B Finetuned
+JSL MedLlama 8B 0.5726 0.8901 0.8303 0.2758 0.0016 0.4478 0.2352 Biomed. Open 8B Pretrained
+Meditron 7B 0.5102 0.8547 - 0.1945 - 0.3691 - Biomed. Closed 7B Finetuned
+BioMistral 7B 0.4402 0.7938 - 0.2117 - 0.1890 - Biomed. Open 7B Instruct
+Mistral 7B v02 0.4847 0.8709 0.8259 0.1989 - 0.4763 0.2355 Gen. Open 7.2B Instruct
+Mistral 7B v03 0.5084 0.8744 - 0.2106 - 0.4824 - Gen. Open 7.2B Instruct
+Gemma2 2B 0.1728 0.8559 0.817 - - - - Gen. Open 2B Instruct
 Table 4: Model accuracy on MedQA vs. Expert MCQ, showing performance difference between the two tasks.
@@
 gpt-3.5-turbo 16 1202 1.331 1 166 0.602 -54.771
-Meta-Llama-3-8B-Instruct 1 166 0.602 0 11 $0$
-Llama3-OpenBioLLM-8B 2 133 1.504 0 5 $0$
-Mixtral-8x7B-Instruct-v0.1 1 235 0.426 0 9 $0$
-claude-3-haiku 0 83 0.000 0 3 $0$
-claude-3-opus 0 79 0.000 0 4 $0$
-gemini-ultra 0 36 0.000 0 5 $0$
-Llama3-OpenBioLLM-70B 3 169 1.775 0 7 $0$
-gemini-pro 0 40 0.000 0 0 $0$
-Meta-Llama-3-70B-Instruct 0 150 0.000 0 8 $0$
-medlm 1 29 3.448 0 2 $0$
+Meta-Llama-3-8B-Instruct 1 166 0.602 0 11 0.000 -
+Llama3-OpenBioLLM-8B 2 133 1.504 0 5 0.000 -
+Mixtral-8x7B-Instruct-v0.1 1 235 0.426 0 9 0.000 -
+claude-3-haiku 0 83 0.000 0 3 0.000 -
+claude-3-opus 0 79 0.000 0 4 0.000 -
+gemini-ultra 0 36 0.000 0 5 0.000 -
+Llama3-OpenBioLLM-70B 3 169 1.775 0 7 0.000 -
+gemini-pro 0 40 0.000 0 0 0.000 -
+Meta-Llama-3-70B-Instruct 0 150 0.000 0 8 0.000 -
+medlm 1 29 3.448 0 2 0.000 -
 Table 19: Clinician blind evaluations of human and LLM answers showing hallucinations, i.e.
@@
 gpt-4-turbo 44 1232 3.571 5 189 2.646 -25.903
-Llama3-OpenBioLLM-70B 6 169 3.550 0 7 $0$
+Llama3-OpenBioLLM-70B 6 169 3.550 0 7 0.000 -
 gpt-3.5-turbo 36 1202 2.995 5 166 3.012 0.568
-gemini-pro 1 40 2.500 0 0 $0$
+gemini-pro 1 40 2.500 0 0 0.000 -
 phi3-mini-128 26 1127 2.307 5 171 2.924 26.745
-claude-3-opus 1 79 1.266 0 4 $0$
-Mixtral-8x7B-Instruct-v0.1 2 235 0.851 0 9 $0$
-claude-3-haiku 0 83 0.000 0 3 $0$
-gemini-ultra 0 36 0.000 0 5 $0$
-medpalm2 0 56 0.000 0 3 $0$
-medlm 0 29 0.000 0 2 $0$
+claude-3-opus 1 79 1.266 0 4 0.000 -
+Mixtral-8x7B-Instruct-v0.1 2 235 0.851 0 9 0.000 -
+claude-3-haiku 0 83 0.000 0 3 0.000 -
+gemini-ultra 0 36 0.000 0 5 0.000 -
+medpalm2 0 56 0.000 0 3 0.000 -
+medlm 0 29 0.000 0 2 0.000 -
 Table 20: Clinician blind evaluations of human and LLM answers showing ommissions, i.e.
@@
 medpalm2 4 56 7.143 1 3 33.333 366.653
-gemini-pro 2 40 5.000 0 0 $0$
-Meta-Llama-3-70B-Instruct 4 150 2.667 0 8 $0$
-Meta-Llama-3-8B-Instruct 3 166 1.807 0 11 $0$
-claude-3-opus 1 79 1.266 0 4 $0$
+gemini-pro 2 40 5.000 0 0 0.000 -
+Meta-Llama-3-70B-Instruct 4 150 2.667 0 8 0.000 -
+Meta-Llama-3-8B-Instruct 3 166 1.807 0 11 0.000 -
+claude-3-opus 1 79 1.266 0 4 0.000 -
 Mixtral-8x7B-Instruct-v0.1 2 235 0.851 1 9 11.111 1205.640
-claude-3-haiku 0 83 0.000 0 3 $0$
-gemini-ultra 0 36 0.000 0 5 $0$
-medlm 0 29 0.000 0 2 $0$
+claude-3-haiku 0 83 0.000 0 3 0.000 -
+gemini-ultra 0 36 0.000 0 5 0.000 -
+medlm 0 29 0.000 0 2 0.000 -
 Table 21: Specialty Human Eval: Clinician ratings for all model responses across specialties showing counts of each specialty, mean of correctness rating, and MCQ counts with percentages where model responses demonstrated the possibility of harm, hallucinations, and omission of relevant information, i.e.